40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Tue Aug 31 15:11:28 2010, Frank, Notes, DAQ, C3:PSL-ACAV_RCTRANSPD broken? 
    Reply  Tue Aug 31 15:46:13 2010, tara, Notes, DAQ, C3:PSL-ACAV_RCTRANSPD broken? 
       Reply  Tue Aug 31 15:52:43 2010, Frank, Notes, DAQ, C3:PSL-ACAV_RCTRANSPD broken? 
          Reply  Tue Aug 31 16:04:30 2010, tara, Notes, DAQ, C3:PSL-ACAV_RCTRANSPD broken? 
             Reply  Tue Aug 31 16:42:04 2010, tara, Notes, DAQ, C3:PSL-ACAV_RCTRANSPD broken? 
Message ID: 331     Entry time: Tue Aug 31 16:42:04 2010     In reply to: 329
Author: tara 
Type: Notes 
Category: DAQ 
Subject: C3:PSL-ACAV_RCTRANSPD broken? 

Quote:

Quote:

was it grey? If yes the cable already connected to the T is the cable going to the DAQ. That's why there is a T already. Don't disconnect any signals without checking in advance what they are and what they are used for. Some of them are required for locking, e.g. the transmitted light is required for the servo as a feedback if it is locked or not for the temp feedback. If you have to disconnect anything do it but put an entry in the elog.

Quote:

Quote:

can you plz check this channel as it is constantly at -14.4797 since yesterday. Did you disconnect it? If yes, plz elog those things!! Everything must be eloged

 I didn't do anything on that channel. I just removed the BNC cable that connect to a T behind the PD and connected it to an oscilloscope this afternoon. I'll take a look.

 

 No, it's the black one that goes to a multimeter. The grey cable to DAQ is untouched.

 I checked the beam. It hit the PD and gives ~500 mV output when only ACAV is locked. when I disconnected the grey cable that connects the PD to DAQ, the

value is still the same (-14.5).

The connection at DAQ looks good.

The cable is good, the signal can go through.

I thought 0the setup in  RCAV.db and ACAV.db files were wrong so I changed it to the red values

                             rcav.db                     acav.db

field DTYP        VMIVME-3113          VMIVME-3123 ->VMIVME-3113

INP                   #c0 S61@                     #c0 S1@  -> #c0 S62@

 ,and now it's working.

I remember changing these before Frank went to Hanford, I might not save the file. Anyway, now I saved it.

The fluctuation in VCOMON goes down to less than +/- 1V.

ELOG V3.1.3-